OpenStudy (anonymous):

A hot air balloon descends at a rate of 320 feet per minute for 3 minutes. Use integers to calculate and represent the balloon's total change in altitude. Note: Remember to show all of the steps that you use to solve the problem. You can use the comments field to explain your work. Your teacher will review each step of your response to ensure you receive proper credit for your answer.

OpenStudy (anonymous):

what is th e initial altitude?

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

Assuming the hot air balloon is high enough so that it won't hit the floor during the descent. We know that every minute the balloon is descending 320 feet right? So in two minutes it will descend 320+320=640 feet. And in 3 minutes it will descend 320+320+320 = 960 feet. So the balloon has descended 960 feet in 3 minutes!

OpenStudy (anonymous):

Another (faster/better) way of doing this is using the formula \[distance = {speed \times time}\] In our case speed = 320 and time = 3.\[320 \times 3 = 960\]

OpenStudy (anonymous):

I have one more question to ask?

OpenStudy (anonymous):

From midnight to 6:00 am, the temperature rose 8 degrees C. At 6:00 am, the temperature was -20 degrees C. What was the temperature at midnight?

OpenStudy (anonymous):

-20+8